# Spectral Masking and Interpolation Attack (SMIA): A Black-box Adversarial Attack against Voice Authentication and Anti-Spoofing Systems

### v1

@@ 6.5. Stealth and Detectability Analysis
To assess the stealth of our attack, we conducted a visual analysis comparing the spectrograms of our adversarial samples against those from Hai et al. (Hai et al., 2023), as visualized in Figure 4. The spectrograms for the Hai et al. attack (Figures 4(b) and 3(g)) reveal a critical vulnerability: the method introduces highly structured and conspicuous artifacts in the form of clear “silent areas”. These predictable, zeroed-out regions are easily detectable by simple forensic analysis, and if filtered out, the attack’s effectiveness would drop dramatically. In contrast, all modes of our SMIA attack produce more subtle and randomly distributed perturbations that blend into the natural texture of the spectrogram. By avoiding obvious, structured patterns, SMIA generates adversarial samples that are inherently stealthier and more resilient to detection and removal by forensic countermeasures.
(a) Original
(b) Hai et al. (Hai et al., 2023)
(c) SMIA–Mask
(d) SMIA–Interpol.
(e) SMIA–Hybrid
(f) Original
(g) Hai et al. (Hai et al., 2023)
(h) SMIA–Mask
(i) SMIA–Interpol.
(j) SMIA–Hybrid
Figure 3. Visual comparison of attack stealth. The spectrograms show an original audio versus adversarial samples generated by SiFDetectCracker (Hai et al.) (Hai et al., 2023) and our three SMIA modes for two different source audios (top and bottom rows). Note the conspicuous “silent areas” introduced by the SiFDetectCracker (Hai et al., 2023) (b, g), which create a detectable pattern. In contrast, the perturbations from all SMIA modes (c, d, e) and (h, i, j) are more subtle and randomly distributed, making them significantly harder to detect.
@@ 7. Ablation Study
7.1. Ablation Study on Attack Modes
Finally, to understand the contribution of our different perturbation strategies, we conducted an ablation study comparing the Masking, Interpolation, and Hybrid modes. As shown in Table 12, the results highlight that the effectiveness of each mode is highly dependent on the target system’s composition. The key takeaway is that the Interpolation mode proved to be the most potent overall, achieving a perfect 100% ASR in multiple configurations and indicating its superior ability to bypass certain defense combinations.
Attack Mode VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-DARTS
Masking X-Vectors 78.3 68.3 93.3
DeepSpeaker 48.3 71.6 46.7
Interpolation X-Vectors 100 100 68.3
DeepSpeaker 81.3 77.8 23.3
Hybrid X-Vectors 73.3 70 71
DeepSpeaker 50 70 50
7.2. Ablation Study on Randomness
To analyze the impact of our stochastic perturbation strategy, we conducted an ablation study comparing two of our proposed operational modes: a deterministic Simple mode and a stochastic Random mode. The results in Table 13 reveal a crucial insight: while both approaches are exceptionally potent, introducing randomness provides a decisive performance advantage against certain advanced defenses. For example, against the DeepSpeaker (Li et al., 2017) and RawGAT-ST (Tak et al., 2021a) combination, the Random mode boosts the ASR from 77.8% to 93.3%, highlighting the strategic value of stochasticity in overcoming robust countermeasures.
Attack Mode VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-DARTS
Simple X-Vectors 100 99.7 100
DeepSpeaker 96.8 77.8 82.7
Random X-Vectors 100 100 78.3
DeepSpeaker 100 93.3 73
7.3. Stealth Analysis of Simple vs. Random
@@ 8. Discussion
Spectral Masking and Interpolation Attack (SMIA) represents a significant advancement in understanding the vulnerabilities of modern voice biometric security systems. The findings carry several critical implications for the field. The practical success of SMIA, achieving up to a 100% ASR against standalone countermeasures and at least 82.7% against combined, layered systems, underscores a fundamental flaw in current defense paradigms. The evaluation across a variety of models, including open-source standards like X-Vectors and commercial systems such as Microsoft Azure SV, demonstrates the broad, real-world applicability of this attack methodology. A key strength of SMIA is its generalization; by successfully utilizing spoofed audio from the 13 different attack types in the ASVspoof 2019 dataset, this work has demonstrated that SMIA does not depend on the artifacts of a single voice synthesis method. This flexibility positions it as a persistent and evolving threat. Furthermore, the attack’s robustness in simulated over-the-air and over-the-line conditions confirms its practicality beyond purely digital environments.
@@ 9. Future Work
First, to improve the attack’s computational efficiency, the current iterative feedback loop could be replaced with a more sophisticated optimization algorithm. Techniques such as Bayesian optimization or other gradient-free methods could search the parameter space for successful perturbations, potentially reducing the number of queries and overall time required to craft an effective adversarial sample. An even more advanced approach involves training a Deep Neural Network (DNN) to act as a perturbation generator. Such a network could take the initial spoofed audio as input and directly output a perturbed sample or the parameters to generate one, replacing the iterative search with a single, rapid forward pass.
@@ 10. Conclusion
Acknowledgement
This research is supported by the Air Force Office of Scientific Research under award number FA2386-23-1-4003.

### v3

@@ 6.5. Stealth and Detectability Analysis
To assess the stealth of our attack, we conducted a visual analysis comparing the spectrograms of our adversarial samples against those from Hai et al. (hai2023sifdetectcracker), as visualized in Figure 3. The spectrograms for the Hai et al. attack (Figures 3(b) and 3(e)) reveal a critical vulnerability: the method introduces highly structured and conspicuous artifacts in the form of clear “silent areas”. These predictable, zeroed-out regions are easily detectable by simple forensic analysis, and if filtered out, the attack’s effectiveness would drop dramatically. In contrast, all modes of our SMIA attack produce more subtle and randomly distributed perturbations that blend into the natural texture of the spectrogram. By avoiding obvious, structured patterns, SMIA generates adversarial samples that are inherently stealthier and more resilient to detection and removal by forensic countermeasures.
@@ 7. Ablation Study
7.1. Ablation Study on Attack Modes
Finally, to understand the contribution of our different perturbation strategies, we conducted an ablation study comparing the Masking, Interpolation, and Hybrid modes. As shown in Table 10, the results highlight that the effectiveness of each mode is highly dependent on the target system’s composition. While the Interpolation mode appears dominant against RawNet2 (rawnet2) and RawGAT-ST(rawgat) (achieving 100% ASR with X-Vectors), it exhibits a critical vulnerability against RawPC-Darts (rawdarts), where its performance collapses to just 5%. Conversely, the Hybrid mode proves to be the most resilient overall. It sacrifices marginal performance on the simpler defenses to achieve a decisive 84.4% ASR against RawPC-Darts (rawdarts), demonstrating that combining strategies is essential for bypassing the most robust defense combinations.
Attack Mode VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-Darts
Masking X-Vectors 87 75 78
DeepSpeaker 76 72 54
Interpolation X-Vectors 100 100 5
DeepSpeaker 97.7 88 2
Hybrid X-Vectors 94.4 71.1 84.4
DeepSpeaker 73.3 72.2 51.1
7.2. Ablation Study on Randomness
To analyze the impact of our stochastic perturbation strategy, we conducted an ablation study comparing two of our proposed operational modes: a deterministic Simple mode and a stochastic Random mode. The results in Table 11 reveal a crucial insight: while the deterministic approach is highly potent against most systems, introducing randomness provides a decisive performance advantage against specific advanced defenses. Notably, against the DeepSpeaker (deepspeaker) and RawPC-DARTS combination, the Simple mode is effectively blocked (14% ASR). However, the Random mode recovers significant capability, boosting the ASR from 14% to 66.5%, highlighting the strategic value of stochasticity in overcoming robust countermeasures.
Attack Mode VAS Model Anti-Spoofing CM
RawNet2 RawGAT-ST RawPC-DARTS
Simple X-Vectors 100 100 100
DeepSpeaker 93.6 94.4 14
Random X-Vectors 100 100 92.5
DeepSpeaker 93.5 93.5 66.5
7.3. Stealth Analysis of Simple vs. Random
@@ 8. Discussion
SMIA represents a significant advancement in understanding the vulnerabilities of modern voice biometric security systems. The findings carry several critical implications for the field. The practical success of SMIA, achieving up to a 100% ASR against standalone countermeasures and at least 82.7% against combined, layered systems, underscores a fundamental flaw in current defense paradigms. The evaluation across a variety of models, including open-source standards like X-Vectors (xvectors) and commercial systems such as Microsoft Azure SV (microsoft_speaker), demonstrates the broad, real-world applicability of this attack methodology. A key strength of SMIA is its generalization; by successfully utilizing spoofed audio from the 13 different attack types in the ASVspoof 2019 dataset, this work has demonstrated that SMIA does not depend on the artifacts of a single voice synthesis method. This flexibility positions it as a persistent and evolving threat.
@@ 9. Future Work
First, to further enhance the attack’s computational efficiency, the optimization process itself could be replaced entirely. Our Bayesian Optimization can be replaced by a DNN to act as a perturbation generator. Such a network could take the initial spoofed audio as input and directly output a perturbed sample or the parameters to generate one, replacing the multi-query optimization search with a single, rapid forward pass.
@@ 10. Conclusion
Ethical Considerations
This research introduces the SMIA to expose critical security gaps in modern VAS and CMs, with the primary objective of advancing defensive security research. The study demonstrates that current security postures are insufficient against adaptive adversarial attacks and highlights the urgent need for dynamic, context-aware defenses. By responsibly revealing these vulnerabilities, this work aims to assist service providers and researchers in developing next-generation security frameworks capable of detecting sophisticated deepfake threats.
All experiments were conducted under strict ethical guidelines to ensure data privacy and experimental safety. The evaluation relied exclusively on publicly available datasets, including ASVSpoof 2019 for benchmarking spoofing countermeasures (Todisco2019ASVspoof) and LibriSpeech for end-to-end system evaluation (panayotov2015librispeech). No personal data was collected from real-world victims. Additionally, assessments involving commercial APIs (e.g., Microsoft Azure (microsoft_speaker)) were performed using these standardized datasets in controlled environments, ensuring no disruption to live services or unauthorized access to legitimate user accounts.
LLMs Usage Considerations
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality.